OpenStudy (anonymous):

Find the y-coordinate of the vertex of the parabola whose equation is y = x2 - x + 2.

OpenStudy (triciaal):

y = x^2- x + 2 equate with y = ax^2 + bx + c x value of vertex =-b/2 a and y value is f(x)